Find the Square matrix M order of two such that M square of 2 - 2M= (-1 0) 2x2
6 3
1
Expert's answer
2020-01-30T08:49:09-0500

Firstly, let x1, x2, x3 and x4 be the corresponding elements of matrix, so that:

"\begin{pmatrix}
 x_1 & x_2 \\
 x_3 & x_4
\end{pmatrix}" * "\begin{pmatrix}
 x_1 & x_2 \\
 x_3 & x_4
\end{pmatrix}" - 2 * "\begin{pmatrix}
 x_1 & x_2 \\
 x_3 & x_4
\end{pmatrix}" = "\begin{pmatrix}
 -1 & 0 \\
 6 & 3
\end{pmatrix}"


When we multiply matrix by itself and multiply it by two as well, we get:

"\begin{pmatrix}
 x_1^2 + x_2x_3 & x_1x_2 + x_2x_4 \\
 x_1x_3 + x_3x_4 & x_2x_3 + x_4^2
\end{pmatrix}" - "\begin{pmatrix}
 2x_1 & 2x_2 \\
 2x_3 & 2x_4
\end{pmatrix}" = "\begin{pmatrix}
 -1 & 0 \\
 6 & 3
\end{pmatrix}"


After matrix subtraction, we get equations (one per each element):

"1) x_1^2 + x_2x_3 - 2x_1 = -1"

"2) x_1x_2 + x_2x_4 - 2x_2 = 0"

"3) x_1x_3 + x_3x_4 - 2x_3 = 6"

"4) x_2x_3 + x_4^2 - 2x_4 = 3"


Let's group the second and the third equations:

"2) x_2(x_1 + x_4 - 2) = 0"

"3) x_3(x_1 + x_4 - 2) = 6"


From the second equation we understand that x2 = 0 or x1 + x4 - 2 = 0. The second assumption is false because if it were true, then the third equation would be:

"3) x_3 * 0 = 6"

which is impossible. Thus, "x_2=0" .


Let's group the first and the fourth equations:

"1) x_2x_3 = -1 + 2x_1 - x_1^2"

"4) x_2x_3 = 3 + 2x_4 - x_4^2"


As x2 = 0, then x2x3 = 0. That means:

"1) x_1^2 - 2x_1 + 1 = 0"

"4) x_4^2 - 2x_4 - 3 = 0"


Let's solve the first equation:

"x_1^2 - 2x_1 + 1 = 0"

"D = b^2 - 4*a*c = 4 - 4 = 0."

That means there is only one solution for this equation:

"x_1 = \frac{-b}{2*a} = \frac{2}{2} = 1"

Thus, "x_1=1".


Let's solve the fourth equation:

"x_4^2 - 2x_4 - 3 = 0"

"D = b^2 - 4*a*c = 4 + 12 = 16; \sqrt{D} = 4;"

"x_{4;1} = \frac{ -b - \sqrt{D}}{2*a} = \frac{2-4}{2} = -1"

"x_{4;2} = \frac{ -b + \sqrt{D}}{2*a} = \frac{2+4}{2} = 3"

So, x4 has two solutions.


Let's get back to the third equation again:

"x_3(x_1 + x_4 - 2) = 6"

"x_3 = \frac{6}{x_1 + x_4 - 2} = \frac{6}{x_4 - 1}"


If x4 = -1, then:

"x_3 = \frac{6}{-2} = -3"


If x4 = 3, then:

"x_3 = \frac{6}{2} = 3"


Thus, (x3;x4) is (-3;-1) or (3;3).


So, there are two solutions, the matrices are as follows:


1) "\begin{pmatrix}
 1 & 0 \\
 -3 &-1
\end{pmatrix}"


2) "\begin{pmatrix}
 1 & 0 \\
 3 &3
\end{pmatrix}"